When 50,000 mL of oxygen at 20.0 °C is compressed to 5.00 liters, what must the new temperature be to maintain constant pressure?
1
Expert's answer
2020-04-06T08:55:27-0400

For processes with constant pressure

"\frac{V_1}{V_2}=\frac{T_1}{T_2}."

Hence, the final temperature

"T_2=T_1\frac{V_2}{V_1}=(20+273)\:\rm K\times \frac{5.00\:\rm L}{50.0\:\rm L}""=29.3\:\rm K=-244\:\rm \degree C"